OpenStudy (anonymous):

solve each system of equations by substitution y=2x+7 3x-y=-9

OpenStudy (anonymous):

3x - y = -9 3x = y-9 Substituting y = 2x+7 in 3x = y-9 3x = 2x+7-9 3x - 2x = -2 x = -2 Substituting x = -2 in y = 2x+7 y = 2(-2) + 7 = 3